# Jump-teaching: Ultra Efficient and Robust Learning with Noisy Label

### v1

@@ 1 Introduction
Figure 1: Test accuracy and "disagreement" of different strategies with symmetric noise ratio $\epsilon=0.5$, $\epsilon=0.8$. We use Intersection over Union (IoU) between two selections to represent the disagreement. In the cross-update strategy, the disagreement emerges in a different network, while in the other two strategies, it emerges during different epochs. More details are discussed in Appendix A.11
@@
Our contributions can be summarized as follows: 1) We exploit the disagreement between intervalic iterations during training for a single network, and leverage the jump-update strategy to suppress the accumulated error. To the best of our knowledge, it is the first time that a single network has surpassed the capacity of a dual network in the robustness and efficiency of LNL. (2) The loss of sample selection is proposed with a lightweight design and easily selects clean data by intrinsic characteristics. 3) Our proposed approach achieves almost up to 2.53×2.53$\times$2.53 × speedup, 0.46×0.46$\times$0.46 × peak memory footprint, and outperforms state-of-the-art methods in terms of robustness with various noise settings, especially for real-time scenarios. Besides, there is considerable flexibility in integrating this method with other LNL approaches.
@@ 3.1 Jump-update Strategy
Strategy Description sTo simplify our discussion, we dive into the procedure of sample selection and give some definitions. Sample selection is accomplished by the execution of model updates and clean sample selection with multiple iterations. In other words, model updates and clean sample selection are performed only once in every iteration. Previous strategies and ours are shown in Fig. 2(2(a)). Suppose $t_{i}$ denotes the current $i$-th iteration, ancestor iteration refers to the former iteration $t_{j},0\leq j\leq{i-2}$, excluding the previous $(i-1)$-th iteration, during the procedure of sample selection. Similarly, descendant iteration denotes the future iteration $t_{s},i+1\leq s\leq N_{iterations}$. $N_{iterations}$ is the total number of training iterations. In our paradigm, the current network in the $i$-th iteration is trained with clean samples selected only by the network from an ancestor iteration $t_{j}$. This behavior of sample selection exhibits a jump form. The origin of the algorithm’s name is derived from this point. Concretely, we leverage a binary identifier to represent the outcome of the label judgment after clean sample selection. Thus, a binary identification table corresponds to the entire data. The jump-update strategy is divided into three steps: 1) The neural network in current interaction $t_{i}$ generates the new binary identification table by the clean sample selection for the descendant iteration $t_{s}$. 2) We fetch and cache the old binary identification table from the ancestor iteration $t_{j}$. 3) The network updates the parameters based on the clean data judged only by the old table. Before the update, the network inherits the weight from the previous iteration $t_{i-1}$. When the jump-update strategy is applied, disagreements between different training iterations of a neural network appear, leading to a decrease in the accumulated error. Suppose $T$ denotes the jump steps, it should be noted that $T=i-j=s-i,2\leq T\leq N$. The evidence of disagreement and the setting of $T$ are illustrated in Sec. 4.2.
@@
Theoretical Analysis Sample selection inevitably has a selection bias, leading to accumulated error in an error flow. As shown in Fig. 2(2(b)), the graph of error flow presents a sequential form by iterations while the jump-update strategy splits a sequential error flow into multiple error sub-flows. Intuitively, the more training iterations increase, the more rapidly the error is accumulated. Nonetheless, the degree of error accumulation in the two is significantly different. In the sequential form, errors are accumulated consecutively. As the error sub-flows are orthogonal in the jump-update strategy, each error is accumulated only in its own error sub-flow. Thus, the jump-update strategy has a significantly smaller degree of accumulated error compared to the sequential form. This is the source of the magic of the jump-update strategy. We formalize the aforementioned procedure mathematically below and give some detail properties.
@@
Conclusion The jump-update strategy is more effective and efficient than previous works. Compared to methods that implicitly leverage Property 1, it can reduce $D_{a}$ to a small number i.e., $0.5e$ at a constant cost. Compared to the cross-update strategy that leverage Property 2, it can not only reduce the degree of accumulated error significantly more but also halve the training cost.
@@ 3.2 Semantic Loss Decomposition
Motivation According to memorization effects, neural networks prioritize learning simpler patterns from data [49]. The small-loss criterion leverages this effect: samples with clean labels are learned first by the network, hence exhibiting smaller losses compared to noisy samples. However, the relative magnitude of losses is determined by comparison with other samples e.g., rank sample losses with Top-k algorithm[12, 17] or modeling loss distribution[21, 29]. To avoid such costly overheads, we come up with an idea: Is it possible to decompose the single loss and build a distribution? Inspired by memorization effects, if the flipped labels share some identical components with original labels, then for flipped labels, the clean components will be learned first and incur smaller losses, while the noisy components will result in larger losses. This property can be utilized to identify whether a label contains noise. Therefore, we design an auxiliary and lightweight head at the last layer of the network only for sample selection. This head transforms the predicted probability and the one-hot encoding into semantic feature embedding and hash coding, respectively, in other spaces. This enriches the representation of the loss function for sample selection.
@@
Hash Coding with Hardamard Matrix Inspired by [45], we utilize the excellent properties of Hadamard matrices to construct mappings for category encoding. A $K$-bit Hadamard matrix can generate $2K$ codewords, each $K$ bits long, with a minimum Hamming distance of $\frac{K}{2}$. For $K$-bit hash codes, we construct a $K\times K$ Hadamard matrix. From this, we select $c$ row vectors as category encodings, each with a Hamming distance of $\frac{K}{2}$. Labels are mapped to the semantic space through this codebook. This mapping can be presented as $H:\tilde{y}_{i}\in\{0,1\}^{C}\rightarrow\tilde{y}_{i}^{\prime}\in\{-1,1\}^{k}$.
Feature Embedding and Loss Function outputs are mapped to feature embeddings in a subspace via an additional three-layer MLP as detection head, which represented as $f:x_{i}\in\mathbb{R}^{n}\rightarrow z_{i}\in\mathbb{R}^{k}$.
@@ Appendix A Appendix / supplemental material
A.1 Proof of Property 1
@@
A.2 Discussion on Error Flow
@@
A.3 Disccusion on Intial Bias
@@
Figure 4: Effect of initial bias on training
A.4 The Summary of the Benchmark Datasets
@@ A.6 The Simulation of Noise
A.7 A Brief Overview of Compared Methods
@@
A.8 The Structures of Backbones and the Auxiliary Head
@@ A.9 Analysis of Threshold
Table 7: Test Accuracies(%) with different threshold settings
@@
A.10 Ablation Study
We evaluate three components of Jump-teaching: data augmentation(AUG), jump-update strategy(JU) and sample selection(SS). Results of ablation study are shown in Table 8. Effectiveness of Data Augmentation When only a limited number of samples are available, such as in the presence of symmetric noise ratio $\epsilon=0.8$, data augmentation can significantly enhance the learning capabilities of these samples, enabling the network to focus more on the intrinsic characteristics of the images themselves. Effectiveness of Jump-update Strategy The jump-update strategy can effectively overcome error accumulation, allowing it to perform excellently with extreme noise. Without the jump-update strategy, the network is almost unable to learn with the symmetric noise ratio $\epsilon=0.8$ condition on CIFAR-10.
Effectiveness of Sample Selection If noisy samples are not filtered out at all, the network can easily fit them, resulting in reduced generalization performance. This has been verified with various noise conditions.
Table 8: Ablation Study
@@ A.10 Ablation Study
A.11 Discussion about Disaggrement
@@
A.12 The Choice of Jump Step.
We test the effect of different jump steps on Jump-teaching with the results shown in Table 9. Performance across various steps remains similar as error accumulation has been effectively constrained to the epoch level. In the Jump-update Strategy, we opt for a step size of 2, obviating the need for additional forward propagation post-training. Furthermore, empirical evidence indicates a marginal decline in performance as step size increases, corroborating the assertions in Sec. 2 that delays in updating tables can detrimentally impact the network’s capacity to assimilate newly introduced clean samples.
Table 9: Test accuracy(%) on CIFAR-10 and CIFAR-100 with symmetric and asymmetric noise with various noise settings
@@
A.13 The Size of Hadamard Codebook.

### v4

@@ 1 Introduction
Figure 1: Test accuracy and "disagreement" of different strategies with symmetric noise ratio $\epsilon=0.5$, $\epsilon=0.8$. We use Intersection over Union (IoU) between two selections to represent the disagreement. In the cross-update strategy, the disagreement emerges in a different network, while in the other two strategies, it emerges during different epochs. More details are discussed in Appendix A.15.
@@
Our contributions can be summarized as follows: 1) We exploit the disagreement between intervalic iterations during training for a single network, and leverage the jump-update strategy to suppress the accumulated error. To the best of our knowledge, it is the first time that a single network outperforms a dual network in terms of robustness and efficiency in LNL. (2) The loss of sample selection is proposed with a lightweight design and easily selects clean data by intrinsic characteristics. 3) Our proposed approach achieves almost up to 2.53×2.53$\times$2.53 × speedup, 0.46×0.46$\times$0.46 × peak memory footprint, and outperforms state-of-the-art methods in terms of robustness with various noise settings, especially for real-time scenarios. Besides, there is considerable flexibility in integrating this method with other LNL approaches.
@@ 3.1 Jump-update Strategy
Strategy Description. To simplify our discussion, we dive into the procedure of sample selection and give some definitions. Sample selection is accomplished by the execution of model updates and clean sample selection with multiple iterations. In other words, model updates and clean sample selection are performed only once in every iteration. Previous strategies and ours are shown in Fig. 2(2(a)). Suppose $t_{i}$ denotes the current $i$-th iteration, ancestor iteration refers to the former iteration $t_{j},0\leq j\leq{i-2}$, excluding the previous $(i-1)$-th iteration, during the procedure of sample selection. Similarly, descendant iteration denotes the future iteration $t_{s},i+1\leq s\leq N_{iterations}$. $N_{iterations}$ is the total number of training iterations. In our paradigm, the current network in the $i$-th iteration is trained with clean samples selected only by the network from an ancestor iteration $t_{j}$. This behavior of sample selection exhibits a jump form. The origin of the algorithm’s name is derived from this point. Concretely, we leverage a binary identifier to represent the outcome of the label judgment after clean sample selection. Thus, a binary identification table corresponds to the entire data. The jump-update strategy is divided into three steps: 1) The neural network in current interaction $t_{i}$ generates the new binary identification table by the clean sample selection for the descendant iteration $t_{s}$. 2) We fetch and cache the old binary identification table from the ancestor iteration $t_{j}$. 3) The network updates the parameters based on the clean data judged only by the old table. Before the update, the network inherits the weight from the previous iteration $t_{i-1}$. When the jump-update strategy is applied, disagreements between different training iterations of a neural network appear, leading to a decrease in the accumulated error. Suppose $T$ denotes the jump steps, it should be noted that $T=i-j=s-i,2\leq T\leq N$. The evidence of disagreement and the setting of $T$ are illustrated in Sec. 4.2.
@@
Empirical Analysis. Sample selection inevitably has a selection bias, leading to accumulated error in an error flow. As shown in Fig. 2(2(b)), the graph of error flow presents a sequential form by iterations while the jump-update strategy splits a sequential error flow into multiple error sub-flows. Intuitively, the more training iterations increase, the more rapidly the error is accumulated. Nonetheless, the degree of error accumulation in the two is significantly different. In the sequential form, errors are accumulated consecutively. As the error sub-flows are orthogonal in the jump-update strategy, each error is accumulated only in its own error sub-flow. Thus, the jump-update strategy has a significantly smaller degree of accumulated error compared to the sequential form. This is the source of the magic of the jump-update strategy. We formalize the aforementioned procedure mathematically below and give some detail properties.
@@
Conclusion. The jump-update strategy is more effective and efficient than previous works. Compared to methods that implicitly leverage Property 1, it can reduce $D_{a}$ to a small number i.e., $0.5e$ at a constant cost. Compared to the cross-update strategy that leverage Property 2, it can not only reduce the degree of accumulated error significantly more but also halve the training cost.
@@ 3.2 Semantic Loss Decomposition
Motivation. According to memorization effects, neural networks prioritize learning simpler patterns from data [49]. The small-loss criterion leverages this effect: samples with clean labels are learned first by the network, hence exhibiting smaller losses compared to noisy samples. However, the relative magnitude of losses is determined by comparison with other samples e.g., rank sample losses with Top-k algorithm[12, 17] or modeling loss distribution[21, 29]. To avoid such costly overheads, we come up with an idea: Is it possible to decompose the single loss and build a distribution? Inspired by memorization effects, if the flipped labels share some identical components with original labels, then for flipped labels, the clean components will be learned first and incur smaller losses, while the noisy components will result in larger losses. This property can be utilized to identify whether a label contains noise. Therefore, we design an auxiliary and lightweight head at the last layer of the network only for sample selection. This head transforms the predicted probability and the one-hot encoding into semantic feature embedding and hash coding, respectively, in other spaces. This enriches the representation of the loss function for sample selection.
@@
Hash Coding with Hardamard Matrix. Inspired by [45], we utilize the excellent properties of Hadamard matrices to construct mappings for category encoding. A $K$-bit Hadamard matrix can generate $2K$ codewords, each $K$ bits long, with a minimum Hamming distance of $\frac{K}{2}$. For $K$-bit hash codes, we construct a $K\times K$ Hadamard matrix. From this, we select $c$ row vectors as category encodings, each with a Hamming distance of $\frac{K}{2}$. Labels are mapped to the semantic space through this codebook. This mapping can be presented as $H:\tilde{y}_{i}\in\{0,1\}^{C}\rightarrow\tilde{y}_{i}^{\prime}\in\{-1,1\}^{k}$.
Feature Embedding and Loss Function. Outputs are mapped to feature embeddings in a subspace via an additional three-layer MLP as detection head, which represented as $f:x_{i}\in\mathbb{R}^{n}\rightarrow z_{i}\in\mathbb{R}^{k}$.
@@ Appendix A Appendix / supplemental material
A.1 Workflow of Jump-teaching
Figure 4: The framework and workflow of Jump-teaching. Motivated by two key issues in the sample selection method, Jump-teaching proposes a novel model-update strategy to mitigate the selection bias and a plug-and-play sample selection component to replace complex modeling.
A.2 Proof of Property 1
@@
A.3 Discussion on Error Flow
@@
A.4 Disccusion on Intial Bias
@@
Figure 5: Effect of initial bias on training.
A.5 Loss Convergence
We present results for four different temperature values to show how the different values of Temperature scaling affect the convergence of the losses. When T=1, the Binary Cross-Entropy (BCE) loss fails to converge properly. The cases for T=2, 3 and 4 illustrate the effects of temperature scaling. Specifically, temperature scaling slows down the convergence speed of the classification head, which can be obviously observed during the warm-up phase, while it accelerates convergence of the detection head which can be obviously observed during the training phase. Notably, as the temperature T increases, the convergence rate of the Cross-Entropy (CE) loss decreases.
(a) Convergence of Loss BCE
(b) Convergence of Loss CE
Figure 6: Convergence of losses with different temperature scaling factors $T$ under CIFAR-100 Sym. $\epsilon=0.8$.
A.6 Details of Sample Selecion
We include the number and ratio of clean samples selected in each epoch. This data is reported for each noise condition in the CIFAR-10 and CIFAR-100 datasets and is presented in Fig. 7. It can be observed that the number of clean samples continuously increases and stabilizes during the last 20 epochs, presented in Fig. 7(a) and Fig. 7(c). As for Fig. 7(b) and Fig. 7(d), the ratio of clean samples increases steadily during the early and middle stages of training, indicating that the network’s selection ability improves as its performance increases. However, in the later stages (epochs range from 150 to 200), the ratio of clean samples begins to decline, primarily due to the model starting to fit the noisy data, which results in more noisy samples being misclassified as clean.
(a) Clean num. on CIFAR-10
(b) Clean rat. on CIFAR-10
(c) Clean num. on CIFAR-100
(d) Clean rat. on CIFAR-100
Figure 7: The clean number and the clean ratio of selected samples on CIFAR-10 and CIFAR-100 with different noise types and ratios.
A.7 More Experiments on Different Noise Types
We have added the experiments of pairflip-45 and instance-dependent label noise i.e. IDN. Results are shown in Tab. 6 and Tab. 7. Coupled with the noises in the original paper, our experiments support up to five types of noise. On pairflip noise, our algorithm also exhibits its strong performance, with an increase of 6.73% and 9.95% on CIFAR-10 and CIFAR-100, respectively.
Our algorithm maintains a huge advantage in instance-dependent noise. The leading margins are 3.95% (CIFAR-10 IDN. 0.2), 6.74% (CIFAR-10 IDN. 0.4), 10.2% (CIFAR-10 IDN. 0.6), 4.87% (CIFAR-100 IDN. 0.2), 4.3% (CIFAR-100 IDN. 0.4) and 12.67% (CIFAR-100 IDN. 0.6). This indicates that our algorithm is also extremely robust under IDN settings.
Table 6: Performance of methods on CIFAR-10 and CIFAR-100 with pairflip $\epsilon=0.45$ noise.
Methods CIFAR-10 CIFAR-100
CE 50.22 $\pm$ 0.43 21.59 $\pm$ 0.87
Co-teaching 52.99 $\pm$ 0.16 33.22 $\pm$ 0.66
Co-teaching+ 55.19 $\pm$ 0.27 29.26 $\pm$ 0.15
PENCIL 54.58 $\pm$ 2.26 24.97 $\pm$ 0.57
SPRL 90.54 $\pm$ 0.02 53.62 $\pm$ 1.07
FINE 77.09 $\pm$ 0.10 41.62 $\pm$ 1.01
Topo 84.60 $\pm$ 0.45 52.40 $\pm$ 1.42
Jump-teaching 91.33 $\pm$ 0.23 62.35 $\pm$ 0.74
Table 7: Performance of methods on CIFAR-10 and CIFAR-100 with varying IDN noise ratios.
Methods CIFAR-10 CIFAR-100
0.2 0.4 0.6 0.2 0.4 0.6
CE 85.45 $\pm$ 0.57 76.23 $\pm$ 1.54 59.75 $\pm$ 1.30 57.79 $\pm$ 1.25 41.15 $\pm$ 0.83 25.68 $\pm$ 1.55
Co-teaching 88.87 $\pm$ 0.24 73.00 $\pm$ 1.24 62.51 $\pm$ 1.98 43.30 $\pm$ 0.39 23.21 $\pm$ 0.57 12.58 $\pm$ 0.51
Co-teaching+ 89.80 $\pm$ 0.28 73.78 $\pm$ 1.39 59.22 $\pm$ 6.34 41.71 $\pm$ 0.78 24.45 $\pm$ 0.71 12.58 $\pm$ 0.51
PENCIL 86.23 $\pm$ 1.23 69.88 $\pm$ 1.39 44.16 $\pm$ 1.58 57.41 $\pm$ 0.69 48.38 $\pm$ 0.52 29.69 $\pm$ 3.03
FINE 90.85 $\pm$ 1.21 84.53 $\pm$ 0.77 61.24 $\pm$ 2.02 60.14 $\pm$ 0.015 43.05 $\pm$ 1.43 28.88 $\pm$ 2.01
Topo 89.66 $\pm$ 0.86 85.95 $\pm$ 1.02 75.28 $\pm$ 0.64 64.41 $\pm$ 0.37 55.00 $\pm$ 0.82 31.61 $\pm$ 1.45
SPRL 87.99 $\pm$ 0.03 86.15 $\pm$ 0.81 75.01 $\pm$ 0.74 67.66 $\pm$ 0.17 64.55 $\pm$ 0.19 47.16 $\pm$ 1.97
Jump-teaching 94.80 $\pm$ 0.28 92.89 $\pm$ 0.33 85.48 $\pm$ 0.41 72.53 $\pm$ 0.21 68.85 $\pm$ 0.27 59.83 $\pm$ 0.09
A.8 The Summary of the Benchmark Datasets
@@ A.10 The Simulation of Noise
A.11 A Brief Overview of Compared Methods
@@
A.12 The Structures of Backbones and the Auxiliary Head
@@ A.13 Analysis of Threshold
Table 9: Test Accuracies(%) with different threshold settings.
@@
A.14 Ablation Study
We evaluate three components of Jump-teaching: data augmentation(AUG), jump-update strategy(JU) and sample selection(SS). Results of ablation study are shown in Table 10.
Effectiveness of Data Augmentation. When only a limited number of samples are available, such as in the presence of symmetric noise ratio $\epsilon=0.8$, data augmentation can significantly enhance the learning capabilities of these samples, enabling the network to focus more on the intrinsic characteristics of the images themselves.
Effectiveness of Jump-update Strategy. Jump-update strategy can effectively overcome error accumulation, allowing it to perform excellently with extreme noise. Without the jump-update strategy, the network is almost unable to learn with the symmetric noise ratio $\epsilon=0.8$ condition on CIFAR-10.
Effectiveness of Sample Selection. If noisy samples are not filtered out at all, the network can easily fit them, resulting in reduced generalization performance. This has been verified with various noises.
Table 10: Ablation Study.
@@ A.14 Ablation Study
A.15 Discussion about Disaggrement
@@
A.16 The Choice of Jump Step
We test the effect of jump steps on Jump-teaching with the results shown in Table 11. Performance across various steps remains similar as error accumulation has been effectively constrained to the epoch level. In the Jump-update Strategy, we opt for a step size of 2, obviating the need for additional forward propagation post-training. Furthermore, empirical evidence indicates a marginal decline in performance as step size increases, corroborating the assertions in Sec. 2 that delays in updating tables can detrimentally impact the network’s capacity to assimilate newly introduced clean samples.
Table 11: Test accuracy(%) on CIFAR-10 and CIFAR-100 with symmetric and asymmetric noise with various noise settings.
@@ A.16 The Choice of Jump Step
A.17 The Size of Hadamard Codebook